Arecibo L-band Feed ArrayArecibo Galaxy Environment Survey (AGES)

AGES A2048 Observations & Results

Last Update:

Date Observer(s) Log file Output Files and Comments Comments
05/04/2007 RM EM wapp.20070405.a2048.0001 Scan: TOGS calibration, 
wapp.20070405.a2048.0003 Scan: AGES1367_p3_093, 
wapp.20070405.a2048.0004 Scan: AGES1367_p3_095, 
wapp.20070405.a2048.0005 Scan: AGES1367_p3_097, 
wapp.20070405.a2048.0007 Scan: AGES1367_p3_099, started ~2 mins late. Note v. long azimuth drive at transit & derotation of ALFA! Quick re-write of the script to stop things going too horribly wrong. Aborted after 18 minutes
wapp.20070405.a2048.0008 Scan: AGES1367_p3_101, 
wapp.20070406.a2048.0000 Scan: AGES1367_p3_062, Note day change in the fits file! No data showing in the on-line display, hopefully this is a display/midnight issue!
A new AGES begins...
07/04/2007 RM wapp.20070407.a2048.0001 Scan: TOGS calibration, Right on the ZA limit - we seem to be scheduled a bit early today!
wapp.20070407.a2048.0003 Scan: TOGS calibration, Another one while waiting for the field to rise.
wapp.20070407.a2048.0005 Scan: TOGS calibration, This 15-minute schedule quantization might just drive someone crazy.
wapp.20070407.a2048.0007 Scan: AGES1367_p3_064, 
wapp.20070407.a2048.0008 Scan: AGES1367_p3_068, Note that 066 has already been done three times in last years run (due to an accidental repetition)
wapp.20070407.a2048.0009 Scan: AGES1367_p3_070, 
wapp.20070407.a2048.0010 Scan: AGES1367_p3_072, Crossed transit a lot quicker, probably because I left loads of time...
wapp.20070407.a2048.0011 Scan: AGES1367_p3_074, 
wapp.20070407.a2048.0012 Scan: AGES1367_p3_076, No midnight crossing tonight.
Note: have checked files and yesterday's data is all good - it was just a quickview problem. RFI at 1350!
08/04/2007 RM wapp.20070408.a2048.0001 Scan: TOGS Calibration, 
wapp.20070408.a2048.0004 Scan: TOGS Calibration, And now I'm popping over to the stores for a long weight...
wapp.20070408.a2048.0006 Scan: AGES1367_p3_078, 
wapp.20070408.a2048.0007 Scan: AGES1367_p3_080, 10 s power dip - quicklook shows flat spectra. Then 35 s of data. Then flat spectra again. Galfa spectrometer also showing (nearly) flat. Scan aborted. Dropped right out of CIMA to restart everything then restarting with scan p3_082. TOGS also seems to be back now.
wapp.20070408.a2048.0011 Scan: AGES1367_p3_082, Taking data again now.
wapp.20070408.a2048.0012 Scan: AGES1367_p3_084, Transit spin okay today.
wapp.20070408.a2048.0013 Scan: AGES1367_p3_086, 
wapp.20070408.a2048.0014 Scan: AGES1367_p3_088, 
RFI @ 1350.
09/04/2007 RM wapp.20070409.a2048.0002 Scan: TOGS Calibration, 
wapp.20070409.a2048.0004 Scan: TOGS Calibration, 
wapp.20070409.a2048.0006 Scan: AGES1367_p3_090, 
wapp.20070409.a2048.0007 Scan: AGES1367_p3_092, 
wapp.20070408.a2049.0008 Scan: AGES1367_p3_094, 
wapp.20070409.a2048.0009 Scan: AGES1367_p3_096, 
wapp.20070409.a2048.0010 Scan: AGES1367_p3_098, 
wapp.20070408.a2048.0011 Scan: AGES1367_p3_100, 
Cleann run tonight. RFI @ 1350. Web down - log entered manually!
10/04/2007 RM wapp.20070410.a2048.0002 Scan: TOGS calibration, 
wapp.20070410.a2048.0004 Scan: AGES1367_p3_099, 
wapp.20070410.a2048.0005 Scan: AGES1367_p3_080, 
Two scans to finish off. All done.